# Nearest Neighbour Score Estimators for Diffusion Generative Models

### v1

@@ 3.2 Nearest Neighbour Proposals
To quickly identify $\mathcal{K}$, we rely on the Gaussian nature of diffusion processes. Since $p_{t}(\mathbf{x}^{(i)}|\mathbf{z})\propto p_{t}(\mathbf{z}|\mathbf{x}^{(i)})$ and $p_{t}(\mathbf{z}|\mathbf{x}^{(i)})$ is Gaussian, the $k$ most probable elements of the posterior are equivalent to the $k$ elements of $\mathcal{D}$ with the smallest $\ell_{2}$ distance to $\mathbf{z}$. Therefore, we transform the difficult problem of finding the $k$ most likely elements of the posterior into the easier problem of identifying the $k$ nearest neighbours of $\mathbf{z}$ in Euclidean space.
To determine $\mathcal{K}$ in practice, we utilize FAISS (Johnson et al., 2019) to perform fast $k$ nearest neighbour search over $\mathcal{D}$. Specifically, we pre-compute an index $\mathcal{I}$ of $\mathcal{D}$. Then, to construct $q_{t}(\mathbf{x}|\mathbf{z})$, we search $\mathcal{I}$ for the $k$ nearest neighbours of $\mathbf{z}$, returning the nearest neighbour set $\mathcal{K}$ and their $\ell_{2}$ distances $d\in\mathbb{R}^{k}$ to $\mathbf{z}$. Helpfully, since FAISS returns the $\ell_{2}$ distance to $\mathbf{z}$ for each $\mathbf{x}^{(i)}\in\mathcal{K}$, we do not need to retrieve the nearest neighbour elements from our dataset to calculate their proposal probability. Algorithm 1 outlines our method for estimating the marginal score.
@@ 4.1 Analysis
Motivated by our goal of reducing the variance of the score estimate, we derive two theoretical bounds on the trace of the covariance of $\hat{\mu}_{\text{KNN}}$ – our posterior mean estimator. We form our bounds using the trace of the covariance of a Monte Carlo estimator $\hat{\mu}_{MC}$ , and another SNIS estimator with $q_{t}(\mathbf{x}|\mathbf{z})=\frac{1}{N}$ which we denote with $\hat{\mu}_{U}$. We note that $\hat{\mu}_{U}$ is asymptotically equivalent to STF (Xu et al., 2023).
@@
[…]
@@
In general, our estimator’s trace-of-covariance is larger than the Monte Carlo estimator trace-of-covariance. The ratio of the two is upper bounded by the factor $\frac{\mathcal{Z}_{q}}{p_{t}(\mathbf{z})}$ – the ratio of the normalizing constants of the proposal and the posterior distributions. This ratio will be exactly unity when $p_{t}(\mathbf{x}^{(i)}|\mathbf{z})$ is fully concentrated on $\mathcal{K}$ (as is true when $t$ is small or $k=N$). The ratio will also approach unity as $p_{t}(\mathbf{x}^{(i)}|\mathbf{z})$ approaches a uniform distribution over $\mathcal{D}$. In any of these cases, the trace-of-covariance of our estimator will approach the trace-of-covariance of multi-sample posterior Monte Carlo.
@@
[…]
@@ 4.2 Empirical Performance
In Figure 3, we plot score and posterior mean estimator performance across a range of noise levels. At each $t$, we evaluate the estimators using ten thousand $\mathbf{z}$ samples, and one hundred estimator evaluations per $\mathbf{z}$. We compare our method against four others: the single-sample posterior estimator typically used in diffusion and consistency training, a $n=256$ multi-sample posterior Monte Carlo estimator, an STF (Xu et al., 2023) estimator with $n=256$, and EDM – a pre-trained near-SoTA diffusion model (Karras et al., 2022). As EDM is deterministic, we do not report its variance. Further, we do not plot its score metrics since errors in the posterior mean estimation for small $t$ are amplified in the score by a factor of $1/t^{4}$. For the multi-sample Monte Carlo estimator, we calculate the posterior by evaluating Equation 1, incurring a full sum over $\mathcal{D}$ per $\mathbf{z}$.
@@ 5.1 Consistency Models
Method FIDnormal-$\downarrow$ ISnormal-$\uparrow$
CD (Song et al., 2023)
3.55 9.48
CT (Song et al., 2023)
8.70 8.49
iCT (Song & Dhariwal, 2023)
2.83 9.54
iCT (Reimplemented)
3.67 9.45
iCT + STF (Xu et al., 2023)
3.79 9.50
iCT + KNN (ours)
3.59 9.49
From Table 1, we observe consistency training with KNN score estimators results in lower FID and higher Inception Score compared to our re-implementation of iCT (Song & Dhariwal, 2023). While we are unable to reproduce the results of (Song & Dhariwal, 2023) outright (training code was not available at the time of writing) we do improve over our re-implementation of iCT. Surprisingly, we find that using the STF score estimator (Xu et al., 2023) results in worse FID than both our method and the iCT baseline. We hypothesize STF’s bias may contribute to this degradation.
@@ 6 Diffusion Sampling
Figure 5 plots the quality of the resulting hybrid samples against $t_{\mathrm{switch}}$. We compare our estimator, STF, and a baseline which initializes EDM sampling using samples from the forward likelihood. Naively, our estimator is able to drive FID to an unprecedented minimum of 0.5. However, this performance is matched by forward process initialization, suggesting that the improved FID is caused by a reduction in sample generalization.
As shown by Section 5 and by Xu et al. (2023), training diffusion generative models with improved score estimates does not degrade generalization. However, when sampling using an estimator which closely matches the exact minimizer of the diffusion training objective (Equation 6), we see that generalization disappears. This suggests that the lauded generalization of diffusion models can be attributed to bias in the network score estimator caused by architecture, optimization procedure or other factors.
@@
In contrast, beyond $t_{\mathrm{switch}}<2$, the FID of STF increases, indicating it is unsuitable for general PF-ODE integration. We hypothesize that this increase is because for small $t$, the STF estimator’s performance is reliant on including $\mathbf{x}^{(i)}$ used to generate $\mathbf{z}$ in the SNIS batch. For general PF-ODE integration where there is no such $\mathbf{x}^{(i)}$, the uniform proposal of STF is unable to produce a suitable score estimate.
@@ 7 Related Work
Importance Sampling for Generative Model Training Multiple methods have used importance sampling to reduce variance in generative model training. Reweighted Wake-Sleep (Bornschein & Bengio, 2015) uses importance sampling to improve variance in the sleep-phase of the wake-sleep algorithm (Hinton et al., 1995). IWAE (Burda et al., 2016) uses importance sampling to tighten the ELBO used to train variational autoencoders (Kingma & Welling, 2014). Similar to this work, Stable Target Fields (Xu et al., 2023) utilizes an SNIS estimator to estimate the marginal score of diffusion processes. However, their work uses a uniform proposal and is focused on improving diffusion models training instead of generally estimating the score function.
@@ Appendix B Extension to Generalized Diffusion Processes
The derivations of $s(t)$ and $\sigma(t)$ is covered in depth in (Karras et al., 2022). We now derive a general KNN score estimator for a generic diffusion processes defined by $s(t)$ and $\sigma(t)$. First, we define a forward likelihood
@@
$=s(t)^{-d}\cdot(2\pi)^{\frac{-d}{2}}\sigma(t)^{-d}\text{exp}\left(\frac{-\left\lVert\frac{\mathbf{z}}{s(t)}-\mathbf{x}^{(i)}\right\rVert_{2}^{2}}{2s\sigma(t)^{2}}\right)$ (50)
@@
$=\frac{\mathbf{x}^{(i)}-\frac{\mathbf{z}}{s(t)}}{\sigma(t)^{2}}.$ (55)
@@
$=\mathop{\mathbb{E}}_{\mathbf{x}^{(i)}\sim p_{t}(\mathbf{x}^{(i)}|\mathbf{z}/s(t))}\left[\frac{\mathbf{x}^{(i)}-\frac{\mathbf{z}}{s(t)}}{\sigma(t)^{2}}\right]$ (57)
$=\frac{\mathop{\mathbb{E}}\left[\mathbf{x}\Big{|}\frac{\mathbf{z}}{s(t)},t\right]-\frac{\mathbf{z}}{s(t)}}{\sigma(t)^{2}}.$ (58)
@@
Because $p_{t}\left(\mathbf{x}^{(i)}\Big{|}\right)\frac{\mathbf{z}}{s(t)}\propto p_{t}\left(\frac{\mathbf{z}}{s(t)}\Big{|}\mathbf{x}^{(i)}\right)$ and $p_{t}\left(\frac{\mathbf{z}}{s(t)}\Big{|}\mathbf{x}^{(i)}\right)$ is Gaussian, the most probable elements of $p_{t}\left(\mathbf{x}^{(i)}\Big{|}\right)\frac{\mathbf{z}}{s(t)}$ are the elements with the smallest distance $\left\lVert\frac{\mathbf{z}}{s(t)}-\mathbf{x}^{(i)}\right\rVert||_{2}^{2}$. We can therefore use a fast KNN search with query $\frac{\mathbf{z}}{s(t)}$ over $\mathcal{D}$ to identify $\mathcal{K}^{\prime}$. Together, an algorithm for estimating the score of a generic diffusion process is outlined in Algorithm 2
@@
  Input: $\mathbf{z}$, dataset $\mathcal{D}$, index $\mathcal{I}$, scale $s(t)$, noise level $\sigma(t)$
  $\mathcal{K},d\leftarrow\texttt{search}(I,\mathbf{z}/s(t))$
  $q_{t}(\mathbf{x}^{(i)}|\frac{\mathbf{z}}{s(t)})\leftarrow\text{exp}(\frac{-d_{i}^{2}}{2\sigma(t)^{2}})\ \forall\mathbf{x}^{(i)}\in\mathcal{K}$Equation 61}
  $q_{t}(\mathbf{x}^{(i)}|\frac{\mathbf{z}}{s(t)})\leftarrow\text{exp}(\frac{-d_{k}^{2}}{2\sigma(t)^{2}})\ \forall\mathbf{x}^{(i)}\notin\mathcal{K}$
  $q_{t}(\mathbf{x}^{(i)}|\frac{\mathbf{z}}{s(t)})\leftarrow\frac{q(\mathbf{x}^{(i)}|\frac{\mathbf{z}}{s(t)})}{\sum_{j=1}^{|\mathcal{K}|}q(\mathbf{x}^{(j)}|\frac{\mathbf{z}}{s(t)})}$
@@
  return $\frac{\hat{\mathbf{x}}-\frac{\mathbf{z}}{s(t)}}{\sigma(t)^{2}}$

### v2

@@ 3.2 Nearest Neighbour Proposals
There are several beneficial properties of our nearest neighbours proposal. Most importantly, Equation 13, approximates the posterior using only the elements of $\mathcal{K}$, instead requiring a full summation over $\mathcal{D}$. In addition, since $p_{t}(\mathbf{z}|\mathbf{x})=\mathcal{N}(\mathbf{z}|\mathbf{x},t^{2}\mathbf{I})$ is Gaussian, and the size of the $\mathcal{D}$ is known, the normalizing constant $\mathcal{Z}_{q}=\sum_{i=1}^{k}p_{t}(\mathbf{z}|\mathbf{x}_{i})+(N-k)p_{t}(\mathbf{z}|\mathbf{x}_{k})$ is found easily. Finally, the shape of $q_{t}(\mathbf{x}^{(i)}|\mathbf{z})$ matches the shape of $p_{t}(\mathbf{x}^{(i)}|\mathbf{z})$, as demonstrated in Figure 1.
Examining Figure 1, for $\mathbf{x}^{(i)}\in\mathcal{K}$, the proposal is proportional to the posterior with $\mathcal{Z}_{q}q_{t}(\mathbf{x}^{(i)}|\mathbf{z})=p_{t}(\mathbf{z})p_{t}(\mathbf{x}^{(i)}|\mathbf{z})$. In general, $p_{t}(\mathbf{z})\leq\mathcal{Z}_{q}$ with equality in two cases. When $p_{t}(\mathbf{z}|\mathbf{x}_{k})=0$ or $k=N$, the posterior is fully concentrated on $\mathcal{K}$ and $p_{t}(\mathbf{z})=\sum_{i=1}^{N}p_{t}(\mathbf{z}|\mathbf{x}^{(i)})=\sum_{i=1}^{k}p_{t}(\mathbf{z}|\mathbf{x}_{i})$. In these cases, our nearest neighbour proposal exactly matches the posterior distribution.
When $p_{t}(\mathbf{z}|\mathbf{x}_{k})>0$ and $k<N$, the likelihood $p_{t}(\mathbf{z}|\mathbf{x}_{k})$ upper bounds the likelihood for all $\mathbf{x}^{(i)}\notin\mathcal{K}$. In this case, we underweight the probability for $\mathbf{x}^{(i)}\in\mathcal{K}$ and assign more mass to the tails of the proposal distribution, as seen in Figure 1b.
Equation 13 requires access to the $k$ most probable elements of the posterior. To quickly identify $\mathcal{K}$, we rely on the Gaussian nature of diffusion processes. Since $p_{t}(\mathbf{x}^{(i)}|\mathbf{z})\propto p_{t}(\mathbf{z}|\mathbf{x}^{(i)})$ and $p_{t}(\mathbf{z}|\mathbf{x}^{(i)})$ is Gaussian, the $k$ most probable elements of the posterior are equivalent to the $k$ elements of $\mathcal{D}$ with the smallest $\ell_{2}$ distance to $\mathbf{z}$. Therefore, we transform the difficult problem of finding the $k$ most likely elements of the posterior into the easier problem of identifying the $k$ nearest neighbours of $\mathbf{z}$ in Euclidean space.
To determine $\mathcal{K}$ in practice, we utilize FAISS (Johnson et al., 2019) to perform fast $k$ nearest neighbour search over $\mathcal{D}$. To construct $q_{t}(\mathbf{x}^{(i)}|\mathbf{z})$ for a given $\mathbf{z}$, we first use FAISS to find the nearest neighbours of $\mathbf{z}$. Helpfully, FAISS returns both the nearest neighbour set $\mathcal{K}$ and their $\ell_{2}$ distances $d\in\mathbb{R}^{k}$ to $\mathbf{z}$. Using $d$, we compute the forward likelihood for each element of $\mathcal{K}$, and use Equation 13 to compute the proposal. We can then use the proposal to estimate the score via Equation 11. Algorithm 1 outlines our full nearest neighbour score estimation algorithm.
@@ 4.1 Analysis
Motivated by our goal of reducing the variance of the score estimate, we derive two theoretical bounds on the trace of the covariance of $\hat{\mu}_{\text{KNN}}$ – our posterior mean estimator. We form our bounds using the trace of the covariance of a Monte Carlo estimator $\hat{\mu}_{MC}$ and another SNIS estimator with $q_{t}(\mathbf{x}^{(i)}|\mathbf{z})=\frac{1}{N}$ which we denote with $\hat{\mu}_{U}$. We note that $\hat{\mu}_{U}$ is is only asymptotically equivalent to STF (Xu et al., 2023) because in addition to a uniform $q_{t}(\mathbf{x}^{(i)}|\mathbf{z})$, STF also deterministically includes the source $\mathbf{x}^{(i)}$ in their sample batch $S$.
@@
Let $t\in(0,\infty)$, $\hat{\mu}_{\text{MC}}$ be the Monte Carlo estimator defined by Equation 9, and $\hat{\mu}_{\text{KNN}}$ be the estimator described by Equation 11 with proposal given by Equation 13. Then for a fixed $n\geq 1$,
@@
In general, our estimator’s trace-of-covariance is larger than the simple Monte Carlo estimator’s trace-of-covariance. The ratio of the two is upper bounded by the factor $\frac{\mathcal{Z}_{q}}{p_{t}(\mathbf{z})}$ – the ratio of the normalizing constants of the proposal and the posterior distributions. This ratio will be exactly unity when $p_{t}(\mathbf{x}^{(i)}|\mathbf{z})$ is fully concentrated on $\mathcal{K}$ (as is true when $t$ is small or $k=N$). The ratio will also approach unity as $p_{t}(\mathbf{x}^{(i)}|\mathbf{z})$ approaches a uniform distribution over $\mathcal{D}$. In any of these cases, the trace-of-covariance of our estimator will approach the trace-of-covariance of multi-sample posterior Monte Carlo.
@@
Let $\hat{\mu}_{\text{MC}}$ be defined by Equation 9. Let $\hat{\mu}_{\text{KNN}}$ and $\hat{\mu}_{\text{U}}$ be two estimators defined by Equation 11 with proposals given by Equation 13 and $q_{t}(\mathbf{x}^{(i)}|\mathbf{z})=\frac{1}{N}$ respectively. Then, for $t\in(0,\infty)$ and a fixed $n\geq 1$,
$\text{Tr}\left(\text{Cov}\left(\hat{\mu}_{\text{KNN}}\right)\right)$$\leq$
@@
$+$ $\frac{(N-k)}{N}\text{Tr}\left(\text{Cov}\left(\hat{\mu}_{\text{U}}\right)\right).$
@@ 4.2 Empirical Performance
In Figure 3, we plot score and posterior mean estimator performance across a range of noise levels. At each $t$, we evaluate the estimators using ten thousand $\mathbf{z}$ samples, and one hundred estimator evaluations per $\mathbf{z}$. We compare our method against four others: the single-sample posterior estimator typically used in diffusion and consistency training, a $n=256$ multi-sample posterior Monte Carlo estimator, an STF (Xu et al., 2023) estimator with $n=256$, and EDM – a pre-trained near-SoTA diffusion model (Karras et al., 2022). As EDM is deterministic, we do not report its variance. Further, we do not plot its score metrics since errors in the posterior mean estimation for small $t$ are amplified in the score by a factor of $1/t^{4}$. For the multi-sample Monte Carlo estimator, we calculate the posterior by calculating Equation 1.
@@ 5.1 Consistency Models
[…]
CD (Song et al., 2023) 3.55 9.48
CT (Song et al., 2023) 8.70 8.49
iCT (Song & Dhariwal, 2024) 2.83 9.54
iCT (Reimplemented) 3.67 9.45
iCT + STF (Xu et al., 2023) 3.79 9.50
iCT + KNN (ours) 3.59 9.49
From Table 1, we observe consistency training with KNN score estimators results in lower FID and higher Inception Score compared to our re-implementation of iCT (Song & Dhariwal, 2024). While we are unable to reproduce the results of (Song & Dhariwal, 2024) outright (training code was not available at the time of writing) we do improve over our re-implementation of iCT. Surprisingly, we find that using the STF score estimator (Xu et al., 2023) results in worse FID than both our method and the iCT baseline. We hypothesize STF’s bias may contribute to this degradation.
@@ 6 Diffusion Sampling
Figure 5 plots the quality of the resulting hybrid samples against $t_{\mathrm{switch}}$. We compare our estimator, STF, and a baseline which initializes EDM sampling using samples from the forward likelihood. Naively, our estimator is able to drive FID to an unprecedented minimum of 0.5. However, this performance is matched by forward process initialization, indicating that integrating the PF-ODE with our estimator is equivalent to drawing samples from the training dataset. This means our method cannot be used on its own as a generative model, as it does not produce generalization.
As shown by Section 5 and by Xu et al. (2023), training diffusion generative models with improved score estimates does not degrade generalization. However, when sampling using an estimator which closely matches the exact minimizer of the diffusion training objective (Equation 6), we see that generalization disappears. This reinforces the findings of (Yi et al., 2023): that the lauded generalization properties of diffusion models can be attributed to bias in the network score estimator caused by architecture, optimization procedure or other factors.
@@
In contrast, when $t_{\mathrm{switch}}<2$, the FID of STF increases, indicating it is unsuitable for general purpose PF-ODE integration. We hypothesize that this increase is because for small $t$, the STF estimator’s performance is reliant on including $\mathbf{x}^{(i)}$ used to generate $\mathbf{z}$ in the SNIS batch. For general PF-ODE integration where there is no such $\mathbf{x}^{(i)}$, the uniform proposal of STF is unable to produce a suitable score estimate.
@@ 7 Related Work
Importance Sampling for Generative Model Training Multiple methods have used importance sampling to reduce variance in generative model training. Reweighted Wake-Sleep (Bornschein & Bengio, 2015) uses importance sampling to improve variance in the sleep-phase of the wake-sleep algorithm (Hinton et al., 1995). IWAE (Burda et al., 2016) uses importance sampling to tighten the ELBO used to train variational autoencoders (Kingma & Welling, 2014). Importance sampling has also been utilized to minimize bias in diffusion training (Kim et al., 2024). Similar to this work, Stable Target Fields (Xu et al., 2023) utilizes an SNIS estimator to estimate the marginal score of diffusion processes. However, their work uses a uniform proposal and is focused on improving diffusion models training instead of generally estimating the score function.
@@ Appendix B Extension to Generalized Diffusion Processes
The derivations of $s(t)$ and $\sigma(t)$ is covered in depth in (Karras et al., 2022). We now derive a general KNN score estimator for a generic diffusion processes defined by $s(t)$ and $\sigma(t)$. First, we define the forward likelihood for the generalized diffusion process.
@@
$=s(t)^{-d}\cdot(2\pi)^{\frac{-d}{2}}\sigma(t)^{-d}\text{exp}\left(\frac{-\left\lVert\frac{\mathbf{z}}{s(t)}-\mathbf{x}^{(i)}\right\rVert_{2}^{2}}{2\sigma(t)^{2}}\right)$ (50)
@@
$=\frac{\mathbf{x}^{(i)}-\frac{\mathbf{z}}{s(t)}}{s(t)\sigma(t)^{2}}.$ (55)
@@
$=\mathop{\mathbb{E}}_{\mathbf{x}^{(i)}\sim p_{t}(\mathbf{x}^{(i)}|\mathbf{z}/s(t))}\left[\frac{\mathbf{x}^{(i)}-\frac{\mathbf{z}}{s(t)}}{s(t)\sigma(t)^{2}}\right]$ (57)
$=\frac{\mathop{\mathbb{E}}\left[\mathbf{x}\Big{|}\frac{\mathbf{z}}{s(t)},t\right]-\frac{\mathbf{z}}{s(t)}}{s(t)\sigma(t)^{2}}.$ (58)
@@
Because $p_{t}\left(\mathbf{x}^{(i)}\Big{|}\frac{\mathbf{z}}{s(t)}\right)\propto p_{t}\left(\frac{\mathbf{z}}{s(t)}\Big{|}\mathbf{x}^{(i)}\right)$ and $p_{t}\left(\frac{\mathbf{z}}{s(t)}\Big{|}\mathbf{x}^{(i)}\right)$ is Gaussian, the most probable elements of $p_{t}\left(\mathbf{x}^{(i)}\Big{|}\frac{\mathbf{z}}{s(t)}\right)$ are the elements with the smallest distance $\left\lVert\frac{\mathbf{z}}{s(t)}-\mathbf{x}^{(i)}\right\rVert_{2}^{2}$. We can therefore use a fast KNN search with query $\frac{\mathbf{z}}{s(t)}$ over $\mathcal{D}$ to identify $\mathcal{K}^{\prime}$. Together, an algorithm for estimating the score of a generic diffusion process is outlined in Algorithm 2
@@
  Input: $\mathbf{z}$, dataset $\mathcal{D}$, index $\mathcal{I}$, neighbour size $k$, sample size $n$, scale $s(t)$, noise level $\sigma(t)$
  $\mathcal{K}^{\prime},d\leftarrow\texttt{search}(I,\mathbf{z}/s(t),k)$
  $p_{t}(\frac{\mathbf{z}}{s(t)}|\mathbf{x}_{i})\leftarrow\text{exp}(\frac{-d_{i}^{2}}{2\sigma(t)^{2}})\ \forall\ \mathbf{x}_{i}\in\mathcal{K}^{\prime}$
  $\mathcal{Z}_{q}\leftarrow\sum_{i=1}^{k}p_{t}(\frac{\mathbf{z}}{s(t)}|\mathbf{x}_{i})+(N-k)p_{t}(\frac{\mathbf{z}}{s(t)}|\mathbf{x}_{k})$
  $q_{t}(\mathbf{x}^{(i)}|\frac{\mathbf{z}}{s(t)})\leftarrow\frac{1}{\mathcal{Z}_{q}}\begin{cases}&p_{t}(\frac{\mathbf{z}}{s(t)}|\mathbf{x}_{i})\ \forall\ \mathbf{x}^{(i)}\in\mathcal{K}^{\prime}\\ &p_{t}(\frac{\mathbf{z}}{s(t)}|\mathbf{x}_{k})\ \forall\ \mathbf{x}^{(i)}\notin\mathcal{K}^{\prime}\end{cases}$
@@
  return $\frac{\hat{\mathbf{x}}-\frac{\mathbf{z}}{s(t)}}{s(t)\sigma(t)^{2}}$